# Agnostic Sharpness-Aware Minimization

### v2

@@ Abstract
Sharpness-aware minimization (SAM) has been instrumental in improving deep neural network training by minimizing both the training loss and the sharpness of the loss landscape, leading the model into flatter minima that are associated with better generalization properties. In another aspect, Model-Agnostic Meta-Learning (MAML) is a framework designed to improve the adaptability of models. MAML optimizes a set of meta-models that are specifically tailored for quick adaptation to multiple tasks with minimal fine-tuning steps and can generalize well with limited data. In this work, we explore the connection between SAM and MAML, particularly in terms of enhancing model generalization. We introduce Agnostic-SAM, a novel approach that combines the principles of both SAM and MAML. Agnostic-SAM adapts the core idea of SAM by optimizing the model towards wider local minima using training data, while concurrently maintaining low loss values on validation data. By doing so, it seeks flatter minima that are not only robust to small perturbations but also less vulnerable to data distributional shift problems. Our experimental results demonstrate that Agnostic-SAM significantly improves generalization over baselines across a range of datasets and under challenging conditions such as noisy labels and data limitation.
@@ 1 Introduction
Deep neural networks have become the preferred method for analyzing data, surpassing traditional machine learning models in complex tasks like classification. These networks process input through numerous parameters and operations to predict classes. Learning involves finding parameters within a model space that minimize errors or maximize performance for a given task. Typically, training data, denoted as $S$, is finite and drawn from an unknown true data distribution $\mathcal{D}$. Larger or more aligned training sets lead to more efficient models.
@@
Several studies have been done on this problem, both theoretically and practically. Statistical learning theory has proposed a number of different complexity measures [37] [7] [27, 8, 32] that are capable of controlling generalization errors. In general, they develop a bound for general error on $\mathcal{D}$. Theory suggests that minimizing the intractable general error on $\mathcal{D}$ is equivalent to minimizing the empirical loss on $S$ with some constraints to the complexity of models and training size [3].
@@
An alternative strategy for mitigating generalization errors involves the utilization of an optimizer to learn optimal parameters for models with a specific local geometry. This approach enables models to locate wider local minima, known as flat minima, which makes them more robust against data shift between train and test sets [20, 30, 13]. The connection between a model’s generalization and the width of minima has been investigated theoretically and empirically in many studies, notably [18, 28, 12, 17]. A specific method within this paradigm is Sharpness-aware Minimisation (SAM) [16], which has emerged as an effective technique for enhancing the generalization ability of deep learning models. SAM seeks a perturbed model within the vicinity of a current model that maximizes the loss over a training set. Eventually, SAM leads the model to the region where both the current model and its perturbation model have low loss values, which ensure flatness. The success of SAM and its variants [24, 23, 36] has inspired further investigation into its formulation and behavior, as evidenced by recent works such as [21, 26, 5].
@@
SAM significantly enhances robustness against shifts between training and test datasets, thereby reducing overfitting and improving overall performance across different datasets and domains. This robust optimization approach aligns particularly well with the principles of Model-Agnostic Meta-Learning (MAML) [14]. MAML aims to find a set of meta-model parameters that not only generalize well on current tasks but can also be quickly adapted to a wide range of new tasks. Furthermore, the agnostic perspective of MAML is particularly enticing for enhancing generalization ability because it endeavors to learn the optimal meta-model from meta-training sets capable of achieving minimal losses on independent meta-testing sets, thus harmonizing with the goal of generalization.
@@
In this paper, inspired by MAML, we initially approach the problem of learning the best model over a training set from an agnostic viewpoint. Subsequently, we harness this perspective with sharpness-aware minimization to formulate an agnostic optimization problem. Additionally, a naive solution akin to MAML does not suit our objectives. Hence, we propose a novel solution for this agnostic optimization problem, resulting in an approach called AgnosticSAM. In summary, our contributions in this work are as follows:
@@
  • We proposed a framework inspired by SAM and MAML works, called Agnostic-SAM to improve model flatness and robustness against noise. Agnostic-SAM update model to a region that minimizes the sharpness of the model on the training set while also implicitly performing well on the validation set by using a combination of gradients on both training and validation sets.
@@
  • We examine standard image classification tasks across multiple datasets and models to demonstrate the effectiveness of Agnostic-SAM in enhancing generalization error compared to baseline methods. Additionally, we assess our approach under noisy label conditions with varying levels of noise. The consistent improvement in performance is evidence that Agnostic-SAM not only enhances robustness against label noise but also contributes to more stable and reliable predictions across different settings.
@@ 2 Related works
The correlation between the wider minima and the generalization capacity has been extensively explored both theoretically and empirically in various studies [20, 30, 13]. Many works suggested that finding flat minimizers might help to reduce generlization error and increase robustness to data distributional shift problems in various settings [20, 30, 13]. There are multiple works have explored the impact of different training parameters, including batch size, learning rate, gradient covariance, and dropout, on the flatness of discovered minima such as [22, 19, 39].
Sharpness-aware minimization (SAM) [16] is a recent optimization technique designed to improve the generalization error of neural networks by considering the sharpness of the loss landscape during training. SAM minimizes the worst-case loss around the current model and effectively updates models towards flatter minima to achieve low training loss and maximize generalization performance on new and unseen data. SAM has been successfully applied to various tasks and domains, such as vision models [11], language models [6], federated learning [33], Bayesian Neural Networks [29], domain generalization [9], multi-task learning [31] and meta-learning bi-level optimization [1]. In  [1], authors discussed SAM’s effectiveness in enhancing meta-learning bi-level optimization, while SAM’s superior convergence rates in federated learning compared to existing approaches in  [33] along with proposing a generalization bound for the global model.
@@
Model-agnostic machine learning techniques have significantly advanced, offering flexible solutions applicable across various models and tasks. Notable methods like LIME [35] and SHAP [25] provide model-agnostic interpretability, allowing users to understand complex models by approximating them with simpler, interpretable models or by attributing predictions to individual features. Additionally, MAML [14] stands out as the most compelling model-agnostic technique that formulates meta-learning as an optimization problem, enabling models to adapt to new tasks with minimal task-specific modifications quickly. Subsequent research has largely focused on addressing the computational challenges of MAML [10, 38] or proposing novel approaches that exploit the concept of model agnostic from MAML across a wide range of tasks, including non-stationary environments [2], alternative optimization strategies [34], and uncertainty estimation for robust adaptation [15].
@@ 3.1 Notions
We start with introducing the notions used throughout our paper. We denote $\mathcal{D}$ as the data/label distribution to generate pairs of data/label $(x,y)$. Given a model with the model parameter $\theta$, we denote the per sample loss induced by $(x,y)$ as $\ell(x,y;\theta)$. Let $S$ be the training set drawn from the distribution $\mathcal{D}$. We denote the empirical and general losses as $\mathcal{L}_{S}\left(\theta\right)=\mathbb{E}_{S}\left[\ell\left(x,y;\theta\right)\right]$ and $\mathcal{L}_{\mathcal{D}}\left(\theta\right)=\mathbb{E}_{\mathcal{D}}\left[\ell\left(x,y;\theta\right)\right]$ respectively. Moreover, we define $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S\right)$ as an upper bound defined over $S$ of the general loss $\mathcal{L}_{\mathcal{D}}\left(\theta\right)$. Finally, we use $|A|$ to denote the cardinality of a set $A$.
@@ 3.2 Problem Formulation
Given a training set $S^{t}$ whose examples are sampled from $\mathcal{D}$ (i.e., $S^{t}\sim\mathcal{D}^{N_{t}}$ with $N_{t}=|S^{t}|$), we use $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{t}\right)$ to train models and among the models that minimize this loss, we select the one that minimize the general loss as follows:
@@
The reason for the formulation in (1) is that although $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S_{t}\right)$ is an upper bound of the general loss $\mathcal{L}_{\mathcal{D}}\left(\theta\right)$, there always exists a gap between them. Therefore, the additional outer minimization helps to refine the solutions. We now denote $S^{v}$ (i.e., $S^{v}\sim\mathcal{D}^{N_{v}}$ with $N_{v}=|S^{v}|$) as a valid set sampled from $\mathcal{D}$. With respect to this valid set, we have the following theorem.
@@
Denote $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S\right):=\max_{\theta^{\prime}:\|\theta^{\prime}-\theta\|_{2}\leq\rho}\mathcal{L}_{S}\left(\theta^{\prime}\right)$. Under some mild condition similar to SAM [16], with a probability greater than $1-\delta$ (i.e., $\delta\in[0,1]$) over the choice of $S^{v}\sim\mathcal{D}^{N_{v}}$, we then have for any optimal models $\theta^{*}\in\mathcal{A}_{\mathcal{D}}(S^{t})$:
@@
Our theorem 1 (proof can be found in Appendix A.1) can be viewed as an extension of Theorem 1 in [16], where we apply the Bayes-PAC theorem from [4] to prove an upper bound for the general loss of any bounded loss, instead of the 0-1 loss in [16]. Moreover, we can generalize this proof for $S^{t}$ to explain why using $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{t}\right):=\max_{\theta^{\prime}:\|\theta^{\prime}-\theta\|_{2}\leq\rho}\mathcal{L}_{S^{t}}\left(\theta^{\prime}\right)$ as an objective to minimize, as in (1). Thanks to Theorem 1, we can rewrite the objectives in (1) as:
@@ 3.3 Our Solution
Our agnostic formulation has the same form as MAML [14], developed for meta-learning. Inspired by MAML, a naive approach is to consider $\theta^{*}=\theta^{*}(\theta)$ and then minimize $\mathcal{L}_{\mathcal{D}}\left({\theta}^{*}\left(\theta\right)\mid S^{v}\right)$ with respect to $\theta$. However, this mainly optimizes the loss $\mathcal{L}_{\mathcal{D}}\left(\theta^{*}\left(\theta\right)\mid S^{v}\right)$ over the validation set $S^{v}$, which is not appropriate for our motivation, where we aim to primarily optimize the loss over the training set $S^{t}$, while using $S^{v}$ to further enhance the generalization ability.
We interpret the bi-level optimization problem in (3) as follows: at each iteration, our primary objective is to optimize $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{t}\right)$, primarily based on its gradients, in such a way that future models are able to implicitly perform well on $S^{v}$. To achieve this, similar to SAM [16], we approximate $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{t}\right)=\max_{\|\theta^{\prime}-\theta\|\leq\rho}\mathcal{L}_{S^{t}}\left(\theta^{\prime}\right)\approx\mathcal{L}_{S^{t}}\left(\theta+\eta_{1}\nabla\mathcal{L}_{S^{t}}\left(\theta\right)\right)$ for a sufficient small learning rate $\eta_{1}>0$ (i.e., $\eta_{1}\|\nabla\mathcal{L}_{S^{t}}\left(\theta\right)\|\leq\rho$) and $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{v}\right)=\max_{\|\theta^{\prime}-\theta\|\leq\rho}\mathcal{L}_{S^{v}}\left(\theta^{\prime}\right)\approx\mathcal{L}_{S^{v}}\left(\theta+\eta_{2}\nabla\mathcal{L}_{S^{v}}\left(\theta\right)\right)$ for a sufficient small learning rate $\eta_{2}>0$ (i.e., $\eta_{2}\|\nabla\mathcal{L}_{S^{v}}\left(\theta\right)\|\leq\rho$).
Moreover, at each iteration, while primarily using the gradients of $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{t}\right)$ for optimization, we also utilize the gradient of $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{v}\right)$ in an auxiliary manner to ensure congruent behavior between these two gradients. Specifically, at the $l$-th iteration, we update as follows:
@@
From (7), we reach the conclusion that the update in (6) (i.e., $\theta_{l+1}=\theta_{l}-\eta\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\tilde{\theta}_{l}^{t}\right)$) aims to minimize simultaneously (i) $\mathcal{\mathcal{L}}_{B_{t}}\left(\theta_{l}\right)$, (ii) $\|\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\|_{2}^{2}$, and maximize (iii) $\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\cdot\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$. While the effects in (i) and (ii) are similar to SAM [16], maximizing $\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\cdot\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right))$ encourages two gradients of the losses over $B^{t}$ and $B^{v}$ to become more congruent.
@@
Theorem 2 (proof can be found in Appendix A.1) indicates that two gradients $\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\tilde{\theta}_{l}^{t}\right)$ and $\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$ are encouraged to be more congruent since our update aims to maximize its lower bound $c\times\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\cdot\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$ (i.e., $c=0.5$ or $c=1.5$). Notice that the negative gradient $-\eta\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\tilde{\theta}_{l}^{t}\right)$ is used to update $\theta_{l}$ to $\theta_{l+1}$, hence this update can have implicit impact on minimizing $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{v}\right)$ since the negative gradient $-\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$ targets to minimize $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{v}\right)=\max_{\|\theta^{\prime}-\theta\|\leq\rho}\mathcal{L}_{S^{v}}\left(\theta^{\prime}\right)\approx\mathcal{L}_{S^{v}}\left(\theta+\eta_{2}\nabla\mathcal{L}_{S^{v}}\left(\theta\right)\right)$.
@@
For the practical Agnostic-SAM algorithm, inspired by SAM [16], we replace $\eta_{1}=\rho_{1}\frac{\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)}{\|\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\|_{2}}$ and $\eta_{2}=\rho_{2}\frac{\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)}{\|\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)\|_{2}}$, where $\rho_{1}>0$ and $\rho_{2}>0$ are perturbation radii. Furthermore, instead of using two fixed sets $S^{t}$ and $S^{v}$, and then sampling mini-batches $B^{t}\sim S^{t}$ and $B^{v}\sim S^{v}$, we employ an on-the-fly strategy. Specifically, let $S$ be the entire training set (e.g., $S=S^{t}\cup S^{v}$), we sample a mini-batch $B\sim S$ and then split it into $B^{t}$ and $B^{v}$ (e.g., $|B^{t}|=0.7|B|$ and $|B^{v}|=0.3|B|$). The pseudo-code of Agnostic-SAM is summarized in Algorithm 1.
@@
Input: $\rho_{1},\rho_{2},\eta$, the number of iterations $L_{\text{iter}}$, and the training set $S$.

### v3

@@ Abstract
Sharpness-aware minimization (SAM) has been instrumental in improving deep neural network training by minimizing both the training loss and the sharpness of the loss landscape, leading the model into flatter minima that are associated with better generalization properties. In another aspect, Model-Agnostic Meta-Learning (MAML) is a framework designed to improve the adaptability of models. MAML optimizes a set of meta-models that are specifically tailored for quick adaptation to multiple tasks with minimal fine-tuning steps and can generalize well with limited data. In this work, we explore the connection between SAM and MAML in enhancing model generalization. We introduce Agnostic-SAM, a novel approach that combines the principles of both SAM and MAML. Agnostic-SAM adapts the core idea of SAM by optimizing the model toward wider local minima using training data, while concurrently maintaining low loss values on validation data. By doing so, it seeks flatter minima that are not only robust to small perturbations but also less vulnerable to data distributional shift problems. Our experimental results demonstrate that Agnostic-SAM significantly improves generalization over baselines across a range of datasets and under challenging conditions such as noisy labels or data limitation.
@@ 1 Introduction
Deep neural networks have become the preferred method for analyzing data, surpassing traditional machine learning models in complex tasks such as classification. These networks process input through numerous parameters and operations to predict classes. The learning process involves finding parameters within a model space that minimize errors or maximize performance for a given task. Typically, training data, denoted as $S$, is finite and drawn from an unknown true data distribution $\mathcal{D}$. Larger or more aligned training sets lead to more efficient models.
@@
Several studies have been done on this problem, both theoretically and practically. Statistical learning theory has proposed different complexity measures that are capable of controlling generalization errors (Vapnik, 1998; Bartlett & Mendelson, 2003; Mukherjee et al., 2002; Bousquet & Elisseeff, 2002; Poggio et al., 2004). In general, they develop a bound for general error on $\mathcal{D}$. Theory suggests that minimizing the intractable general error on $\mathcal{D}$ is equivalent to minimizing the empirical loss on $S$ with some constraints to the complexity of models and training size (Alquier et al., 2016b). An alternative strategy for mitigating generalization errors involves the utilization of an optimizer to learn optimal parameters for models with a specific local geometry. This approach enables models to locate wider local minima, known as flat minima, which makes them more robust against data shift between training and testing sets (Jiang et al., 2020; Petzka et al., 2021; Dziugaite & Roy, 2017).
@@
The connection between a model’s generalization and the width of minima has been investigated theoretically and empirically in many studies, notably (Hochreiter & Schmidhuber, 1994; Neyshabur et al., 2017; Dinh et al., 2017; Fort & Ganguli, 2019). A specific method within this paradigm is Sharpness-aware Minimisation (SAM) (Foret et al., 2021), which has emerged as an effective technique for enhancing the generalization ability of deep learning models. SAM seeks a perturbed model within the vicinity of a current model that maximizes the loss over a training set. Eventually, SAM leads the model to the region where both the current model and its perturbation model have low loss values, which ensure flatness. The success of SAM and its variants (Kwon et al., 2021; Kim et al., 2022; Truong et al., 2023) has inspired further investigation into its formulation and behavior, as evidenced by recent works such as (Kaddour et al., 2022; Möllenhoff & Khan, 2022; Andriushchenko & Flammarion, 2022).
@@
SAM significantly enhances robustness against shifts between training and testing datasets, thereby reducing overfitting and improving overall performance across different datasets and domains. This robust optimization approach aligns particularly well with the principles of Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017). MAML aims to find a set of meta-model parameters that not only generalize well on current tasks but can also be quickly adapted to a wide range of new tasks. Furthermore, the agnostic perspective of MAML is particularly enticing for enhancing generalization ability because it endeavors to learn the optimal meta-model from meta-training sets capable of achieving minimal losses on independent meta-testing sets, thus harmonizing with the goal of generalization.
@@
In this paper, inspired by MAML and leveraging SAM, we initially approach the problem of learning the best model over a training set from an agnostic viewpoint. Subsequently, we harness this perspective with sharpness-aware minimization to formulate an agnostic optimization problem. However, a naive solution akin to MAML does not suit our objectives. We propose a novel solution for this agnostic optimization problem, resulting in an approach called AgnosticSAM. In summary, our contributions to this work are as follows:
@@
  • We proposed a framework inspired by SAM and MAML works, called Agnostic-SAM to improve model flatness and robustness against noise. Agnostic-SAM updates a model to a region that minimizes the sharpness on the training set while also implicitly performing well on the validation set by using a combination of gradients on both training and validation sets.
@@
  • We demonstrate the effectiveness of Agnostic-SAM in improving generalization performance. Our initial examination focuses on image classification tasks, including training from scratch and transfer learning across a range of datasets, from small to large scale. We also extend this experiment under noisy label conditions with varying levels of noise. Additionally, we apply Agnostic-SAM in MAML settings to validate the effectiveness of our method in generalizing beyond the meta-training tasks and its adaptability across different domains. The consistent improvement in performance across experiments indicates that Agnostic-SAM not only enhances robustness against label noise and improves the model’s generalization across diverse tasks, but also contributes to more stable and reliable predictions in different settings.
@@ 2 Related works
The correlation between the wider minima and the generalization capacity has been extensively explored both theoretically and empirically in various studies (Jiang et al., 2020; Petzka et al., 2021; Dziugaite & Roy, 2017). Many works suggested that finding flat minimizers might help to reduce generlization error and increase robustness to data distributional shift problems in various settings (Jiang et al., 2020; Petzka et al., 2021; Dziugaite & Roy, 2017). There are multiple works have explored the impact of different training parameters, including batch size, learning rate, gradient covariance, and dropout, on the flatness of discovered minima such as (Keskar et al., 2017; Jastrzebski et al., 2017; Wei et al., 2020).
Sharpness-aware minimization (SAM) (Foret et al., 2021) is a recent optimization technique designed to improve the generalization error of neural networks by considering the sharpness of the loss landscape during training. SAM minimizes the worst-case loss around the current model and effectively updates models towards flatter minima to achieve low training loss and maximize generalization performance on new and unseen data. SAM has been successfully applied to various tasks and domains, such as vision models (Chen et al., 2021), language models (Bahri et al., 2022), federated learning (Qu et al., 2022), Bayesian Neural Networks (Nguyen et al., 2023), domain generalization (Cha et al., 2021), multi-task learning (Phan et al., 2022) and meta-learning bi-level optimization (Abbas et al., 2022). In  Abbas et al. (2022), authors discussed SAM’s effectiveness in enhancing meta-learning bi-level optimization, while SAM’s superior convergence rates in federated learning compared to existing approaches in  Qu et al. (2022) along with proposing a generalization bound for the global model. Additionally, multiple varieties of SAM have been proposed (Kwon et al., 2021), (Li et al., 2024), (Du et al., 2022) to tackle the different problems of the original method.
@@
Model-agnostic machine learning techniques have significant advances and offer flexible solutions applicable across various models and tasks. In which, MAML (Finn et al., 2017) stands out as the most compelling model-agnostic technique that formulates meta-learning as an optimization problem, enabling models to improve the model ability to quickly adapt to new tasks with minimal task-specific modifications or limited additional data. Subsequent research has largely focused on addressing the computational challenges of MAML (Chen et al., 2023; Wang et al., 2023) or proposing novel approaches that exploit the concept of model agnostic from MAML across a wide range of tasks, including non-stationary environments (Al-Shedivat et al., 2018), alternative optimization strategies (Rajeswaran et al., 2019), and uncertainty estimation for robust adaptation (Finn et al., 2018). Recently, Abbas et al. (2022) analyzed the loss-landscape of MAML models and proposed the integration of SAM in training to improve the generalization performance of a meta-model.
@@ 3.1 Notions
We start by introducing the notions used throughout our paper. We denote $\mathcal{D}$ as the data/label distribution to generate pairs of data/label $(x,y)$. Given a model with the model parameter $\theta$, we denote the per sample loss induced by $(x,y)$ as $\ell(x,y;\theta)$. Let $S$ be the training set drawn from the distribution $\mathcal{D}$. We denote the empirical and general losses as $\mathcal{L}_{S}\left(\theta\right)=\mathbb{E}_{S}\left[\ell\left(x,y;\theta\right)\right]$ and $\mathcal{L}_{\mathcal{D}}\left(\theta\right)=\mathbb{E}_{\mathcal{D}}\left[\ell\left(x,y;\theta\right)\right]$ respectively. We define $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S\right)$ as an upper bound defined over $S$ of the general loss $\mathcal{L}_{\mathcal{D}}\left(\theta\right)$. Note that inspired by SAM (Foret et al., 2021), we use the sharpness over $S$ to define $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S\right)$. Finally, we use $|A|$ to denote the cardinality of a set $A$.
@@ 3.2 Problem Formulation
Given a training set $S^{t}$ whose examples are sampled from $\mathcal{D}$ (i.e., $S^{t}\sim\mathcal{D}^{N_{t}}$ with $N_{t}=|S^{t}|$), we use $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{t}\right)$ to train models. Among the models that minimize this loss, we select the one that minimizes the general loss as follows:
@@
The reason for the formulation in (1) is that although $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{t}\right)$ is an upper bound of the general loss $\mathcal{L}_{\mathcal{D}}\left(\theta\right)$, there always exists a gap between them. Therefore, the additional outer minimization helps to refine the solutions. We now denote $S^{v}$ (i.e., $S^{v}\sim\mathcal{D}^{N_{v}}$ with $N_{v}=|S^{v}|$) as a valid set sampled from $\mathcal{D}$. With respect to this valid set, we have the following theorem.
@@
Denote $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S\right):=\max_{\theta^{\prime}:\|\theta^{\prime}-\theta\|_{2}\leq\rho}\mathcal{L}_{S}\left(\theta^{\prime}\right)$. Under some mild condition similar to SAM (Foret et al., 2021), with a probability greater than $1-\delta$ (i.e., $\delta\in[0,1]$) over the choice of $S^{v}\sim\mathcal{D}^{N_{v}}$, we then have for any optimal models $\theta^{*}\in\mathcal{A}_{\mathcal{D}}(S^{t})$:
@@
Our theorem 1 (proof can be found in Appendix A.1) can be viewed as an extension of Theorem 1 in Foret et al. (2021), where we apply the Bayes-PAC theorem from Alquier et al. (2016a) to prove an upper bound for the general loss of any bounded loss, instead of the 0-1 loss in Foret et al. (2021). We can generalize this proof for $S^{t}$ to explain why we use $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{t}\right):=\max_{\theta^{\prime}:\|\theta^{\prime}-\theta\|_{2}\leq\rho}\mathcal{L}_{S^{t}}\left(\theta^{\prime}\right)$ as an objective to minimize, as in (1). Based on Theorem 1, we can rewrite the objectives in (1) as:
@@ 3.3 Our Solution
Our motivation here is to primarily optimize the loss over the training set $S^{t}$, while using $S^{v}$ to further enhance the generalization ability. Our agnostic formulation has the same form as MAML  (Finn et al., 2017), developed for meta-learning. Inspired by MAML, a naive approach would be to consider $\theta^{*}=\theta^{*}(\theta)$ and then minimize $\mathcal{L}_{\mathcal{D}}\left({\theta}^{*}\left(\theta\right)\mid S^{v}\right)$ with respect to $\theta$. However, this naive approach does not align with our objective, as it mainly focuses on optimizing the loss $\mathcal{L}_{\mathcal{D}}\left(\theta^{*}\left(\theta\right)\mid S^{v}\right)$ over the validation set $S^{v}$.
We interpret the bi-level optimization problem in (3) as follows: at each iteration, our primary objective is to optimize $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{t}\right)$, primarily based on its gradients, in such a way that future models are able to implicitly perform well on $S^{v}$. To achieve this, similar to SAM (Foret et al., 2021), we approximate $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{t}\right)=\max_{\|\theta^{\prime}-\theta\|\leq\rho}\mathcal{L}_{S^{t}}\left(\theta^{\prime}\right)\approx\mathcal{L}_{S^{t}}\left(\theta+\eta_{1}\nabla\mathcal{L}_{S^{t}}\left(\theta\right)\right)$ for a sufficient small learning rate $\eta_{1}>0$ (i.e., $\eta_{1}\|\nabla\mathcal{L}_{S^{t}}\left(\theta\right)\|\leq\rho$) and $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{v}\right)=\max_{\|\theta^{\prime}-\theta\|\leq\rho}\mathcal{L}_{S^{v}}\left(\theta^{\prime}\right)\approx\mathcal{L}_{S^{v}}\left(\theta+\eta_{2}\nabla\mathcal{L}_{S^{v}}\left(\theta\right)\right)$ for a sufficient small learning rate $\eta_{2}>0$ (i.e., $\eta_{2}\|\nabla\mathcal{L}_{S^{v}}\left(\theta\right)\|\leq\rho$).
At each iteration, while primarily using the gradients of $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{t}\right)$ for optimization, we also utilize the gradient of $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{v}\right)$ in an auxiliary manner to ensure congruent behavior between these two gradients. Specifically, at the $l$-th iteration, we update as follows:
@@
From (7), we reach the conclusion that the update in (6) (i.e., $\theta_{l+1}=\theta_{l}-\eta\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\tilde{\theta}_{l}^{t}\right)$) aims to minimize simultaneously (i) $\mathcal{\mathcal{L}}_{B_{t}}\left(\theta_{l}\right)$, (ii) $\|\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\|_{2}^{2}$, and maximize (iii) $\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\cdot\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$. While the effects in (i) and (ii) are similar to SAM (Foret et al., 2021), maximizing $\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\cdot\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right))$ encourages two gradients of the losses over $B^{t}$ and $B^{v}$ to become more congruent.
@@
Theorem 2 (proof can be found in Appendix A.1) indicates that two gradients $\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\tilde{\theta}_{l}^{t}\right)$ and $\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$ are encouraged to be more congruent since our update aims to maximize its lower bound $c\times\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\cdot\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$ (i.e., $c=0.5$ or $c=1.5$). Notice that the negative gradient $-\eta\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\tilde{\theta}_{l}^{t}\right)$ is used to update $\theta_{l}$ to $\theta_{l+1}$, hence this update can have an implicit impact on minimizing $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{v}\right)$ since the negative gradient $-\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$ targets to minimize $\mathcal{L}_{\mathcal{D}}\left(\theta\mid S^{v}\right)=\max_{\|\theta^{\prime}-\theta\|\leq\rho}\mathcal{L}_{S^{v}}\left(\theta^{\prime}\right)\approx\mathcal{L}_{S^{v}}\left(\theta+\eta_{2}\nabla\mathcal{L}_{S^{v}}\left(\theta\right)\right)$.
@@
Inspired by SAM Foret et al. (2021), we set $\eta_{1}=\rho_{1}\frac{\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)}{\|\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\|_{2}}$ and $\eta_{2}=\rho_{2}\frac{\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\theta_{l}\right)}{\|\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\theta_{l}\right)\|_{2}}$, where $\rho_{1}>0$ and $\rho_{2}>0$ are perturbation radius. Furthermore, instead of splitting the training set $S$ into two fixed subsets, $S^{t}$ and $S^{v}$, which reduces the number of training samples, we set $S^{t}=S^{v}=S$, allowing the entire training set to be used for updating the model. This approach is especially beneficial for training on small datasets. Optionally, we apply momentum with a factor $\beta$ to approximate the gradient of the full validation set using gradients from mini-batches. The effectiveness of this term will be discussed in section 5.
The pseudo-code of Agnostic-SAM is summarized in Algorithm 1.
@@
Input: $\rho_{1},\rho_{2},\eta,\beta$, the number of iterations $L_{\text{iter}}$, and the training set $S$. Initialize gradient on the validation set $g_{v}\leftarrow 0$